OpenStudy (anonymous):

HELP NOW PLEASE!!! Use the Laws of Exponents to simplify the expression three to the power of fifty over three to the power of twenty Explain how you found your answer and write the answer as an exponent.

OpenStudy (anonymous):

3 to the power of 30 is the answer. the law of exponents is that when you devide 2 exponential numbers, subtract the exponents.

OpenStudy (anonymous):

3 ^50 / 3 ^20 = 3 ^(50-20)= 3 ^30, contact me , clasesparticularesss@gmail.com